6. What is the frequency of a microwave whose wavelength is 1.5 x 10 m
Answer #1

here,

the wavelength of microwave , lamda = 1.5 * 10^-2 m

the speed of light , c = 3 * 10^8 m/s

the frequency of wave , f = c/lamda

f = 3 * 10^8 /( 1.5 * 10^-2) Hz

f = 2 * 10^10 Hz

the frequency of wave is 2 * 10^10 Hz